# Feature Request Analysis and Processing: Tasks, Techniques, and Trends

### v1

@@ 7.1. Summarization
This paper presents a systematic survey focused on the analysis and processing of user feature requests, acknowledging their critical role as a direct source of user needs for contemporary software evolution. Employing a defined research methodology that included systematic literature searching, rigorous screening protocols, and snowballing techniques (detailed in Section 3), we identified and thoroughly analyzed 131 relevant primary studies published up to August 2025.
@@ 7.2. Challenges and Opportunities
While this systematic mapping study provides a snapshot of the current landscape, the dynamic nature of software development and user feedback necessitates continuous investigation. We intend to actively monitor the evolving research in feature request analysis and processing to reflect new advancements and shifting trends. Beyond ongoing monitoring, several specific avenues represent critical next steps for the research community.
@@ 9. Conclusion
User feature requests, sourced directly from end-users, are increasingly vital for software evolution but pose unique challenges due to their volume, informality, and divergence from traditional requirements. This paper presents a comprehensive survey of 131 academic articles, mapping the current research landscape for analyzing and processing these requests throughout the software lifecycle. We examine techniques for twelve research topics on feature requests, organizing findings within a requirements engineering framework. The survey concludes by highlighting significant research gaps and future directions.

### v2

@@ 7.1. Summarization
This paper presents a systematic survey focused on the analysis and processing of user feature requests, acknowledging their critical role as a direct source of user needs for contemporary software evolution. Employing a defined research methodology that included systematic literature searching, rigorous screening protocols, and snowballing techniques (detailed in Section 3), we identified and thoroughly analyzed 140 relevant primary studies published up to August 2025.
@@ 7.2. Challenges and Opportunities
While this systematic literature review provides a snapshot of the current landscape, the dynamic nature of software development and user feedback necessitates continuous investigation. We intend to actively monitor the evolving research in feature request analysis and processing to reflect new advancements and shifting trends. Beyond ongoing monitoring, several specific avenues represent critical next steps for the research community.
@@ 9. Conclusion
User feature requests, sourced directly from end-users, are increasingly vital for software evolution but pose unique challenges due to their volume, informality, and divergence from traditional requirements. This paper presents a comprehensive survey of 140 studies, synthesizing the current research landscape on analyzing and processing these requests throughout the software lifecycle. We examine techniques for twelve research topics on feature requests, organizing findings within a requirements engineering framework. The survey concludes by highlighting significant research gaps and future directions.